42: The difference between the square of one even number and the square of the next smaller even number is 52. What is the larger number?
Found 2 solutions by dolly, stanbon:
Answer by dolly(163)

Let one even number be x
so the next even number = x+2
So (x+2)^2 - x^2 = 52
==> x^2 + 4x + 4 - x^2 = 52
==> 4x + 4 = 52
==> 4x = 52 - 4
==> 4x = 48
==> x = 48/4
==> x = 12
So the larger even number = x+2
= 14

Answer by stanbon(75887)

The difference between the square of one even number and the square of the next smaller even number is 52. What is the larger number ?
--------
Let the larger even number be 2(n+1)
Let the smaller even number be 2n
-------------
EQUATION:
(2n+2)^2 - (2n)^2 = 52
(2n)^2+8n+4 - (2n)^2 = 52
8n+4 = 52
8n=48
n=6
2n=12 (the smaller even number)
2n+2=14 (the larger even number)
=============
Cheers,
Stan H.